If light travels for 7.2 minutes, how many meters does it travel?
1
Expert's answer
2019-04-19T10:28:54-0400

We can find the distance traveled by the light from the formula:


"s = ct,"

here, "s" is the distance traveled by the light during time "t", "c = 3 \cdot 10^8 m/s" is the speed of light.

Then, we get:


"s = ct = 3 \cdot 10^8 \dfrac{m}{s} \cdot 7.2 min \cdot \dfrac{60 s}{1 min} = 1.3 \cdot 10^{11} m."

Answer:

"s = 1.3 \cdot 10^{11} m."